kekeman:

Circle A is shown with a central angle marked 30 degrees and the radius marked 5 inches. Which of the following could be used to calculate the area of the sector in the circle shown above? π(5in)30 over 360 π(5in)230 over 360 π(30in)25 over 360 π(30in)5 over 360

Florisalreadytaken:

im pretty sure ive answered this same question at some point we're asked to find the area of just a part of a circle. \[A_p=\frac{\theta}{360} * \pi r^{2}\] we're given that: \[ \theta=30^o \] \[ r=5(in) \] that said, we get this equation: \[ A_p=\frac{30^o}{360} * \pi \times 5(in)^{2} \]

Florisalreadytaken:

or, by changing the places of the variables, to fit your answer it is: \[ A_p= \pi \times 5(in)^{2} \times \frac{30^o}{360} \] so, which option is the right one?

TyDaGreatest:

I never know math

kekeman:

π(5in)230 over 360

Florisalreadytaken:

@florisalreadytaken wrote:
or, by changing the places of the variables, to fit your answer it is: [color= inherit;][/color]\[ A_p= \pi \times 5(in)^{2} \times \frac{30^o}{360} \] so, which option is the right one?
🤦‍♂️🤦‍♂️🤦‍♂️ just look at this

kekeman:

π(5in)^2 30 over 360

kekeman:

Idk why but people keep saying its π(30in)^2 5 over 360

kekeman:

So i am stuck between B.) and C.)

Florisalreadytaken:

oh my lord -- how about you consider looking at A) and D)?

Florisalreadytaken:

@kekeman wrote:
π(5in)^2 30 over 360
and yes, this would be correct

SmokeyBrown:

Here's another way you could think of it: A full circle has 360 degrees, so the section of the circle with 30 degrees will make up 30/360 of the circle, or 1/12 of the full circle. In other words, the π(5in)^2 part represents the area of the entire circle, and the 30/360 part represents the fraction of the entire circle you're interested in. tldr, Florisalreadytaken is right, no cap